OpenStudy (anonymous):

p=b+a+4c solve for b

OpenStudy (dan815):

rearrange for b

OpenStudy (dan815):

subtract a and 4c from both sides

OpenStudy (dan815):

p-a-4c=b+a+4c-a-4c =b